OpenStudy (anonymous):

You earn money to buy DVDs by mowing lawns for x dollars per lawn. One weekend you mow 4 lawns and buy 2 DVDs. Each DVD costs y dollars. You have $8 left at the end of the weekend. Which equation models this situation in slope-intercept form?

OpenStudy (anonymous):

A. -2y=-4x+10 B. x=1/2y+2 C. y=2x+4 D. y=2x-4

OpenStudy (anonymous):

D: help

OpenStudy (callisto):

Perhaps it's like this: One weekend you mow 4 lawns(x) and buy 2 DVDs. Each DVD costs y dollars. You have $8 left: 4x -2y=8 2y = 4x-8 Divide both sides by 2 to get the slope-intercept form y = 2x-4

OpenStudy (anonymous):

oh, that is what i got, but i was confused.

OpenStudy (callisto):

Yes?

OpenStudy (anonymous):

hmm?

OpenStudy (callisto):

What were you confused with?

OpenStudy (anonymous):

I wasn't sure if that was the right answer,

OpenStudy (anonymous):

but thanks!

OpenStudy (callisto):

welcome!~